# Self-Supervised Contrastive Learning is Approximately Supervised Contrastive Learning

### v1

@@ 3.3 Characterizing the representations learned with $\mathcal{L}^{\mathrm{NSCL}}$
Because minimizing over all neural networks $f$ is intractable, we follow prior work (mixon2020neuralcollapseunconstrainedfeatures, ; doi:10.1073/pnas.2103091118, ) and adopt the unconstrained features model. Specifically, we treat $f$ as an arbitrary function that given an augmented sample $x^{l}_{i}$ selects an embedding $z^{l}_{i}$ as a free learnable vector in $\mathbb{R}^{d}$. The following theorem characterizes the global minimizers of the NSCL loss, showing that they satisfy the NC1, NC2, and NC4 properties of NC (doi:10.1073/pnas.2015509117, ; han2022neural, ), and thus learn representations similar to those of minimizers of other classification losses, such as cross-entropy (LU2022224, ; 10.5555/3692070.3692468, ; pmlr-v139-graf21a, ), MSE (han2022neural, ; tirer2022extendedunconstrainedfeaturesmodel, ; zhou2022optimizationlandscapeneuralcollapse, ), and SCL (pmlr-v139-graf21a, ).
@@ 4.1 Experimental Setup
Datasets.  We experiment with the following datasets - CIFAR10 and CIFAR100 (krizhevsky2009learning, ), mini-ImageNet (NIPS2016_90e13578, ), and SVHN (37648, ). For additional experiments and details, see Appendix A.
Methods, architectures and optimizers.  We trained our models with the SimCLR (pmlr-v119-chen20j, ) algorithm. We use a ResNet-50 (He_2016_CVPR, ) encoder with a width-multiplier factor of 2. The projection head follows a standard two-layer MLP architecture composed of: Linear($2048\rightarrow 2048$) $\rightarrow$ ReLU $\rightarrow$ Linear($2048\rightarrow 128$). For additional experiments with MoCo v2 (chen2020improvedbaselinesmomentumcontrastive, ), see Appendix A.
Instead of training our models with the standard InfoNCE loss that is used in SimCLR, we use the DCL loss that avoids positive-negative coupling during training (10.1007/978-3-031-19809-0_38, ). In order to minimize the loss, we adopt the LARS optimizer (you2017large, ) which has been shown in (pmlr-v119-chen20j, ) to be effective for training with large batch sizes. For LARS, we set the momentum to $0.9$ and the weight decay to $1\mathrm{e}^{-6}$. All experiments are carried out with a batch size of $B=1024$. The base learning rate is scaled with batch size as $0.3\cdot\lfloor B/256\rfloor$, following standard practice (pmlr-v119-chen20j, ). We employ a warm-up phase (goyal2017accurate, ) for the first 10 epochs, followed by a cosine learning rate schedule without restarts  (loshchilov2016sgdr, ) for the remaining epochs. All models were trained on a single node with two 94 GB NVIDIA H100 GPUs.
@@
To evaluate the quality of learned representations, we use two methods: the Nearest Class-Center Classification (NCCC) accuracy (galanti2023comparative, ), and linear probing accuracy (kohn2015s, ; gupta2015distributional, ; belinkov2022probing, ). Suppose we have a feature map $f$, a classification task with $C$ classes, and a dataset $S=\cup^{C}_{i=1}S_{i}$ (either training or test data). We estimate $\text{err}^{\text{LP}}_{m,S}(f)$ and $\text{err}^{\text{NCC}}_{m,S}(f)$ with $h_{f,\widehat{S}}$ being an NCC classifier or a linear classifier. To train the linear classifier, we use cross-entropy minimization for 500 epochs, with batch size $\min(|\widehat{S}|,256)$, learning rate $3\mathrm{e}^{-4}$, weight decay $5\mathrm{e}^{-4}$, and momentum 0.9. The expectation over the selection of $\widehat{S}=\cup^{C}_{i=1}\widehat{S}_{i}$ is estimated by averaging the error over 5 selections of $\widehat{S}$ from $S$.
@@ 4.2 Experimental Results
In Fig. 2(top) we observe that the DCL loss consistently upper bounds the NSCL loss, and that the two losses become closer for tasks with a larger number of classes (e.g., CIFAR100 and mini-ImageNet compared to CIFAR10 and SVHN). In addition, when $C$ is large (e.g., $C=100$), the bound becomes very tight, closely matching the losses. Although we notice a slight divergence between the self-supervised and supervised loss metrics as training progresses, the overall trend continues to align well with the inequality stated in Thm. 1. For similar experiments with MoCo-v2, see Fig. 7.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) SVHN (b) CIFAR10 (c) CIFAR100 (d) mini-ImageNet
Figure 2: (Top) We train the model to minimize the DCL loss, tracking the DCL loss, the NSCL loss, and the bound NSCL+$\log(1+\tfrac{n_{\max}\mathrm{e}^{2}}{N-n_{\max}})$ on both the training and test sets throughout training. All three quantities are highly correlated. (Bottom) We compare the NSCL loss of two models: one trained with the DCL loss and the other with the NSCL loss. The resulting NSCL losses are comparable, regardless of the training objective. In both the top and bottom plots, correlations are computed between the DCL and NSCL losses on the train and test data.
@@
For each value of $C$, we randomly sample $C$ classes from the full dataset, train a self-supervised model from scratch on data from these classes only, and compute $\mathcal{L}^{\mathrm{DCL}}(f)-\mathcal{L}^{\mathrm{NSCL}}(f)$ at various training epochs. To account for variability in class selection, we report the averaged value of this quantity over five independent random selections of $C$ classes along with error bars. Fig. 3 presents the empirical results, showing that the gap $\mathcal{L}^{\mathrm{DCL}}(f)-\mathcal{L}^{\mathrm{NSCL}}(f)$ slightly increases over the course of training but remains consistently bounded by $\log(1+\tfrac{\mathrm{e}^{2}}{C-1})$ at all epochs. Moreover, the magnitude of this gap decreases with $C$ and is highly correlated with our bound.
Refer to caption Refer to caption Refer to caption
(a) CIFAR10 (b) CIFAR100 (c) mini-ImageNet
@@
Fig. 4(bottom) compares the bound from Cor. 1 with the NCCC and linear-probe errors of DCL-trained models on 2-way downstream tasks (averaged over 10 tasks), for both train and test data. Unlike Prop. 7 in galanti2023generalizationboundsfewshottransfer , and despite the constants in our bound, it is fairly tight as $m$ increases; for instance, at $m=500$, it indicates that the few-shot errors are below 0.45.At $m=10^{6}$, the test-time bound (red horizontal line) predicts an error of at most 0.25. Fig. 5 shows that DCL training only modestly lowers CDNV but reduces directional CDNV by about an order of magnitude during training. Because Prop. 1 links few-shot error primarily to the directional CDNV, this sharp drop explains why DCL-trained models can recover labels (see Fig. 4) with limited supervision. In contrast, NSCL reduces both variances substantially, which accounts for its lower $m$-shot error.
@@
UMAP visualizations of clustering behaviors.  We visualize the clustering behavior of DCL and NSCL-trained representations learned on mini-ImageNet using 2D UMAP (2018arXivUMAP, ) projections. We randomly selected 5 classes from mini-ImageNet and sampled 200 images per class, projecting their corresponding embeddings into 2D space using UMAP. As shown in Fig. 1, training with the NSCL loss results in tight and clearly separable clusters, as predicted in Thm. 2. Although DCL is label-agnostic, as shown in Thm. 1 and Fig. 2, DCL training implicitly minimizes the NSCL loss, which explains the visible clustering behavior seen in Fig. 1.
@@ 5 Discussion, Limitations, and Future Work
Nevertheless, our work has several important limitations that remain open for future investigation. While we show that the DCL and NSCL losses are close in value, this does not directly imply that their respective minimizers are close in parameter space or yield similar representations. It would be interesting to explore this connection both theoretically and empirically—for example, by monitoring the distance between two models trained using DCL and NSCL on the same batches. Another limitation is that the bound in Prop. 1 (and Cor. 1) involves large constants and scales linearly with $C^{\prime}$, which may limit its practical utility. Tightening this bound by relaxing these dependencies would enhance its predictive power. Finally, extending these ideas to multimodal models such as CLIP (pmlr-v139-radford21a, ), or to LLMs trained with analogous loss functions, presents an exciting direction for future research.

### v2

@@ 3.3 Characterizing the representations learned with $\mathcal{L}^{\mathrm{NSCL}}$
Because minimizing over all neural networks $f$ is intractable, we follow prior work (mixon2020neuralcollapseunconstrainedfeatures; doi:10.1073/pnas.2103091118) and adopt the unconstrained features model. Specifically, we treat $f$ as an arbitrary function that given an augmented sample $x^{l}_{i}$ selects an embedding $z^{l}_{i}$ as a free learnable vector in $\mathbb{R}^{d}$. The following theorem characterizes the global minimizers of the NSCL loss, showing that they satisfy the NC1, NC2, and NC4 properties of NC (doi:10.1073/pnas.2015509117; han2022neural), and thus learn representations similar to those of minimizers of other classification losses, such as cross-entropy (LU2022224; 10.5555/3692070.3692468; pmlr-v139-graf21a), MSE (han2022neural; tirer2022extendedunconstrainedfeaturesmodel; zhou2022optimizationlandscapeneuralcollapse), and SCL (pmlr-v139-graf21a).
@@ 4.1 Experimental Setup
Datasets.  We experiment with the following datasets - CIFAR10 and CIFAR100 (krizhevsky2009learning), mini-ImageNet (NIPS2016_90e13578), Tiny-ImageNet (deep-learning-thu-2020), SVHN (37648) and ImageNet-1K (5206848).
Methods, architectures and optimizers.  We trained our models with the SimCLR (pmlr-v119-chen20j) algorithm. We use a ResNet-50 (He_2016_CVPR) encoder with a width-multiplier factor of 2 and ViT-Base (dosovitskiy2021imageworth16x16words). Both are followed by a projection head with a standard two-layer MLP architecture composed of: Linear($2048\rightarrow 2048$) $\rightarrow$ ReLU $\rightarrow$ Linear($2048\rightarrow 128$). For additional experiments and setup details about Vision Transformers (dosovitskiy2021imageworth16x16words) and MoCo v2 (chen2020improvedbaselinesmomentumcontrastive), see Appendix A.
Instead of training our models with the standard InfoNCE loss that is used in SimCLR, we use the DCL loss that avoids positive-negative coupling during training (10.1007/978-3-031-19809-0_38). In order to minimize the loss, we adopt the LARS optimizer (you2017large) which has been shown in (pmlr-v119-chen20j) to be effective for training with large batch sizes. For LARS, we set the momentum to $0.9$ and the weight decay to $1\mathrm{e}^{-6}$. All experiments are carried out with a batch size of $B=1024$. The base learning rate is scaled with batch size as $0.3\cdot\lfloor B/256\rfloor$, following standard practice (pmlr-v119-chen20j). We employ a warm-up phase (goyal2017accurate) for the first 10 epochs, followed by a cosine learning rate schedule without restarts  (loshchilov2016sgdr) for the remaining epochs. All models were trained on a single node with two 94 GB NVIDIA H100 GPUs.
@@
To evaluate the quality of learned representations, we use two methods: the Nearest Class-Center Classification (NCCC) accuracy (galanti2023comparative), and linear probing accuracy (kohn2015s; gupta2015distributional; belinkov2022probing). Suppose we have a feature map $f$, a classification task with $C$ classes, and a dataset $S=\cup^{C}_{i=1}S_{i}$ (either training or test data). We estimate $\text{err}^{\text{LP}}_{m,S}(f)$ and $\text{err}^{\text{NCC}}_{m,S}(f)$ with $h_{f,\widehat{S}}$ being an NCC classifier or a linear classifier. To train the linear classifier, we use cross-entropy minimization for 500 epochs, with batch size $\min(|\widehat{S}|,256)$, learning rate $3\mathrm{e}^{-4}$, weight decay $5\mathrm{e}^{-4}$, and momentum 0.9. The expectation over the selection of $\widehat{S}=\cup^{C}_{i=1}\widehat{S}_{i}$ is estimated by averaging the error over 5 selections of $\widehat{S}$ from $S$.
@@ 4.2 Experimental Results
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) SVHN (b) CIFAR10 (c) CIFAR100 (d) mini-ImageNet
[…]
Refer to caption Refer to caption Refer to caption
(a) CIFAR10 (b) CIFAR100 (c) mini-ImageNet
@@
In Fig. 3(top) we observe that the DCL loss consistently upper bounds the NSCL loss, and that the two losses become closer for tasks with a larger number of classes (e.g., CIFAR100 and mini-ImageNet compared to CIFAR10 and SVHN). In addition, when $C$ is large (e.g., $C=100$), the bound becomes very tight, closely matching the losses. Although we notice a slight divergence between the self-supervised and supervised loss metrics as training progresses, the overall trend continues to align well with the inequality stated in Thm. 1. For similar experiments with SimCLR-ViT and MoCo-v2, see Fig. 9 and Fig. 10.
@@
For each value of $C$, we randomly sample $C$ classes from the full dataset, train a self-supervised model from scratch on data from these classes only, and compute $\mathcal{L}^{\mathrm{DCL}}(f)-\mathcal{L}^{\mathrm{NSCL}}(f)$ at various training epochs. To account for variability in class selection, we report the averaged value of this quantity over five independent random selections of $C$ classes along with error bars. Fig. 4 presents the empirical results, showing that the gap $\mathcal{L}^{\mathrm{DCL}}(f)-\mathcal{L}^{\mathrm{NSCL}}(f)$ slightly increases over the course of training. As shown in (5), the loss gap is given by $\log\left(1+\tfrac{\sum_{j\neq i,y_{j}=y_{i}}\texttt{exp}(\texttt{sim}(z_{i},z_{j}))}{\sum_{y_{j}\neq y_{i}}\texttt{exp}(\texttt{sim}(z_{i},z_{j}))}\right)$. At initialization, the embeddings $z_{i}$ are randomly distributed in the representation space. During training, the embeddings cluster with respect to their classes thereby increasing $\texttt{sim}(z_{i},z_{j})$ for $y_{j}=y_{i}$ while reducing it for $y_{j}\neq y_{i}$. This leads to a gradual increase in the loss gap as observed in Fig. 4, but it remains consistently bounded by $\log(1+\tfrac{\mathrm{e}^{2}}{C-1})$ at all epochs. Moreover, the magnitude of this gap decreases with $C$ and is highly correlated with our bound.
@@
Fig. 6 (bottom) compares the bound from Cor. 1 with the NCCC and linear‑probe errors of DCL‑trained models on 2‑way downstream tasks (averaged over 10 tasks), for both train and test data. Unlike Prop. 7 in galanti2023generalizationboundsfewshottransfer, and despite the constants in our bound, it is fairly tight as $m$ increases; for instance, at $m=500$ for CIFAR10, it indicates that the few-shot errors are below 0.89. At $m=10^{6}$, the test-time bound (red horizontal line) predicts an error of at most 0.26.
To complement our experiments, we take a publicly available SimCLR-ResNet-50 model pretrained on IM-1K [^1] https://github.com/AndrewAtanov/simclr-pytorch, and perform 2-way $m$-shot NCCC evaluation similar to Fig. 6 (bottom). In Fig. 5, we verify our error bound with pretrained weights on mini-ImageNet and IM-1K.
[Uncaptioned image] [Uncaptioned image]
(a) mini-ImageNet (b) IM-1K
Figure 5: The bound in Cor. 1 is fairly tight for ImageNet pre-trained models. We reproduced Fig. 6 with a ResNet-50 pretrained on IM-1K.
Fig. 7 shows that DCL training only modestly lowers CDNV but reduces directional CDNV by about an order of magnitude during training. Because Prop. 1 links few‑shot error primarily to the directional CDNV, this sharp drop explains why DCL-trained models can recover labels (see Fig. 6) with limited supervision. In contrast, NSCL reduces both variances substantially, which accounts for its lower mm‑shot error.
Augmentation collapse.  We verified augmentation collapse by training DCL ViT models for 2k epochs and measuring cosine similarity (0 indicates no collapse; 1 indicates perfect collapse). For each dataset, we computed the mean cosine similarity between augmentations of the same sample and between augmentations of different samples, using five augmentations per pair across the full dataset.
Dataset CIFAR10 CIFAR100 mini-ImageNet
same diff same diff same diff
DCL 0.92 0.13 0.91 0.11 0.89 0.07
Table 2: Cosine similarity between embeddings of two augmentations of the same sample (“positives”) versus different samples (“negatives”) for a ViT trained with DCL. Positives exhibit markedly higher similarity than negatives.
As shown in Tab. 2, same-sample augmentations achieve a mean cosine similarity of approximately $0.90$, whereas different-sample pairs remain below $0.15$. This consistent gap across CIFAR-10/100 and mini-ImageNet rules out trivial global collapse and indicates that DCL promotes strong augmentation invariance while preserving inter-sample separability.
@@
Representation Alignment.  Thm. 1 shows that the DCL and NSCL objectives have similar values, but this does not guarantee that the embeddings learned by minimizing the two objectives will have identical geometry. To better understand the structure of representations learned by models trained with DCL and NSCL, we run experiments that quantify representation similarity using Centered Kernel Alignment (CKA) (pmlr-v97-kornblith19a) and Representation Similarity Analysis (RSA) (kriegeskorte2008rsa). When embeddings are highly aligned, RSA and CKA are close to 1; with no alignment, they approach 0.
Encoder CIFAR10 CIFAR100 mini-ImageNet
RSA CKA RSA CKA RSA CKA
ResNet-50 0.83 0.81 0.91 0.91 0.87 0.90
ViT-Base 0.86 0.82 0.91 0.91 0.89 0.89
Table 3: DCL and NSCL models trained for 300 epochs with matched initialization and minibatch order show very high representation similarity.
We train DCL and NSCL models with the exact same initialization, same mini-batches, and same hyperparameter configurations as described in Sec. 4.1, for 300 epochs. As shown in Tab. 3, the models achieve very high RSA and CKA values that confirm alignment at the representation level.
@@ 5 Discussion, Limitations, and Future Work
Nevertheless, our work has several important limitations that remain open for future investigation. While we show that the DCL and NSCL losses are close in value, this does not directly imply that their respective minimizers are close in parameter space or yield similar representations. It would be interesting to explore this connection both theoretically and empirically—for example, by monitoring the distance between two models trained using DCL and NSCL on the same batches. Another limitation is that the bound in Prop. 1 (and Cor. 1) involves large constants and scales linearly with $C^{\prime}$, which may limit its practical utility. Tightening this bound by relaxing these dependencies would enhance its predictive power. Finally, extending these ideas to multimodal models such as CLIP (pmlr-v139-radford21a), or to LLMs trained with analogous loss functions, presents an exciting direction for future research.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract outlines both the theoretical and empirical contributions, as well as the main assumptions made in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a limitations section in the main text.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper provides the full set of assumptions for each theoretical result, along with complete and correct proofs in the appendix. All key logical steps in the derivations are clearly presented.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe the experiments and their setup in detail in Sec. 4.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide well-documented code in the supplemental material to reproduce our experimental results. We use publicly available datasets in our work.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, we provide full details required to reproduce our experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We include error bars when appropriate and feasible, given our computational budget. For instance, in Tab. 1 and Fig. 4, where the experiments were averaged over different sample selections or class subsets, we report the results using means and standard deviations or error bars. Due to limited computational resources, we could not afford to repeat the more computationally intensive experiments. Instead, we prioritized running them on multiple datasets.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, we provide information about our compute resources, and memory.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our work conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This is a theoretical paper on self-supervised learning. While it analyzes algorithms that significantly impact widely used applications, the contributions of this work have no direct societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: As a theoretical paper, it does not pose risks of this kind.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We use publicly available datasets and code with appropriate attribution and in accordance with their respective licenses. Specifically:
    • Datasets: CIFAR-10 and CIFAR100 [krizhevsky2009learning], mini-ImageNet [NIPS2016_90e13578] and SVHN [37648] that are explained in Appendix A.
    • Frameworks: Our models are implemented in PyTorch [paszke2017automatic], which is released under the BSD 3 License.
    • Codebases: We build on official or open-source implementations of SimCLR and MoCo, distributed under Apache 2.0 and CC-BY-NC 4.0 Licenses, respectively.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not create any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We use LLMs for editing (e.g., grammar, spelling, word choice).
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.